OpenStudy (anonymous):

The variables x and y vary directly, and x=3 and y=-4. 1. Write an equation that relates to x and y. 2. Find y when x = 2.

OpenStudy (anonymous):

1. y=ax -4=3a =>a=-4/3 y=(-4/3)*x 2. y=(-4/3)*2=-8/3

OpenStudy (anonymous):

Many thanks! ^^

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

And if it was to be x=9 and y=2 instead, would the equation look like this? y=(2/9)*x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

kk thanks